# Towards AI Search Paradigm

### v1

@@ 6.3.4 Prefill-Decode Separating Deployment
The LLM inference process can be divided into two distinct phases: the prefill phase, which processes the input prompt and is computationally intensive (compute-bound), and the decode phase, which generates subsequent tokens autoregressively and is memory bandwidth-intensive (memory-bound). Recognizing these differing characteristics, mature LLM inference frameworks often implement separate deployment and resource allocation for these phases. For instance, systems like Mooncake (Qin et al., 2024), or as implied by the architecture of models like DeepSeek-R1 (DeepSeek-AI et al., 2025), dynamically adjust the ratio of servers or resources dedicated to prefill and decode tasks based on user load and request patterns. This dynamic allocation ensures optimal utilization of computational resources, minimizing bottlenecks and improving throughput. For example, a system might allocate a certain number of nodes/GPUs for parallel prompt processing (prefill) and a different set for token generation (decode), scaling each pool independently.
@@ 6.3.5 Speculative Decoding
Speculative decoding represents a significant advancement in accelerating large language model inference through parallel processing, and existing methods can be categorized into generation and refinement strategies. Generation approaches include predefined fill tokens, which use random initialization but require multiple refinement iterations; retrieval-based methods like REST (He et al., 2023) that leverage exact suffix matching from datastores; and N-gram techniques such as ANPD (Ou et al., 2024) that adapt predictions based on contextual patterns.
Auto-regressive generation represents the most common approach, split between independent and dependent drafters. Independent drafters like SpecDec (Xia et al., 2023) use smaller models to generate tokens verified by larger target models, with improvements like BiLD (Kim et al., 2024) adding confidence thresholds for verification. Dependent drafters reduce computation by sharing resources between draft and target models. Layer skipping methods like Draft&Verify (Zhang et al., 2023a) selectively skip intermediate layers, while dependent head approaches like EAGLE (Li et al., 2024a) add lightweight prediction heads to the target model’s hidden states. Multi-token prediction techniques generate multiple tokens simultaneously. Medusa (Cai et al., 2024) introduced parameter-efficient lightweight decoding heads on pre-trained models, while Amphista (Li et al., 2024c) enhanced this with bi-directional self-attention to model inter-token relationships.
For refinement strategies, single-pass verification dominates the field. Linear verification methods sequentially validate draft tokens against the target model’s distributions, with Fast Inference (Leviathan et al., 2023) introducing speculative sampling to improve acceptance rates. Tree-based verification approaches like SpecInfer (Miao et al., 2023) construct trees of possible completions for parallel exploration, with optimizations like Sequoia’s (Chen et al., 2024) hardware-aware tree optimizer maximizing computational efficiency. Iterative decoding methods permit multiple refinement iterations, with approaches like Jacobi (Teng et al., 2024) reframing auto-regressive generation as an iterative optimization problem that can produce identical output to traditional decoding under greedy sampling.
@@ 7.1 Human Evaluation
where #$\#\mathrm{Win}$ (or #$\#\mathrm{Lose}$) is the number of results generated from the AI search system better (or worse) than the legacy system, and #$\#\mathrm{Tie}$ indicates two results are equally good or bad. In the manual evaluation, we consider the influence of query complexity on the results for the evaluation. Therefore, we sample three categories of simple, moderately complex, and complex queries to comprehensively compare the two systems. Specifically, the test dataset is constructed in advance by sampling real user queries through two distinct selection strategies as follows:
@@ 7.2 Online A/B Test
To evaluate the real-world effectiveness of the AI search system, we deploy it at Baidu Search and conduct an A/B Test comparing it with the legacy search system. When deploying the AI search system at Baidu Search, specific user-side metrics serve as critical indicators of the effectiveness of search systems in enhancing users’ search satisfaction. In the online A/B test, we conduct the experiments with 1.00% real-world web traffic of Baidu search, focusing on metrics directly related to user engagement and reporting the improvements of the proposed AI search system over the online legacy system. As shown in Table 2, compared to the online legacy system, the AI search system achieves the following relative improvements: the change query rate (CQR) decreases by 1.45%; the number of page views (PV) increases by 1.04%; the number of daily active users (DAU) increases by 1.85%; dwell time (Dwell Time) increases by 0.52%. All reported values are statistically significant with $p<$0.05. These results suggest that the proposed AI search system substantially enhances user engagement within the search engine context.

### v2

@@ 6.3.4 Prefill-Decode Separating Deployment
The LLM inference process can be divided into two distinct phases: the prefill phase, which processes the input prompt and is computationally intensive (compute-bound), and the decode phase, which generates subsequent tokens autoregressively and is memory bandwidth-intensive (memory-bound). Recognizing these differing characteristics, mature LLM inference frameworks often implement separate deployment and resource allocation for these phases. For instance, systems like Mooncake [qin2024mooncakekvcachecentricdisaggregatedarchitecture], or as implied by the architecture of models like DeepSeek-R1 [deepseekai2025deepseekr1incentivizingreasoningcapability], dynamically adjust the ratio of servers or resources dedicated to prefill and decode tasks based on user load and request patterns. This dynamic allocation ensures optimal utilization of computational resources, minimizing bottlenecks and improving throughput. For example, a system might allocate a certain number of nodes/GPUs for parallel prompt processing (prefill) and a different set for token generation (decode), scaling each pool independently.
@@ 6.3.5 Speculative Decoding
Speculative decoding represents a significant advancement in accelerating large language model inference through parallel processing, and existing methods can be categorized into generation and refinement strategies. Generation approaches include predefined fill tokens, which use random initialization but require multiple refinement iterations; retrieval-based methods like REST [he2023rest] that leverage exact suffix matching from datastores; and N-gram techniques such as ANPD [ou2024lossless] that adapt predictions based on contextual patterns.
Auto-regressive generation represents the most common approach, split between independent and dependent drafters. Independent drafters like SpecDec [xia2023speculative] use smaller models to generate tokens verified by larger target models, with improvements like BiLD [kim2024speculative] adding confidence thresholds for verification. Dependent drafters reduce computation by sharing resources between draft and target models. Layer skipping methods like Draft&Verify [zhang2023draft] selectively skip intermediate layers, while dependent head approaches like EAGLE [li2024eagle] add lightweight prediction heads to the target model’s hidden states. Multi-token prediction techniques generate multiple tokens simultaneously. Medusa [cai2024medusa] introduced parameter-efficient lightweight decoding heads on pre-trained models, while Amphista [li2024amphista] enhanced this with bi-directional self-attention to model inter-token relationships.
For refinement strategies, single-pass verification dominates the field. Linear verification methods sequentially validate draft tokens against the target model’s distributions, with Fast Inference [leviathan2023fast] introducing speculative sampling to improve acceptance rates. Tree-based verification approaches like SpecInfer [miao2023specinfer] construct trees of possible completions for parallel exploration, with optimizations like Sequoia’s [chen2024sequoia] hardware-aware tree optimizer maximizing computational efficiency. Iterative decoding methods permit multiple refinement iterations, with approaches like Jacobi [teng2024accelerating] reframing auto-regressive generation as an iterative optimization problem that can produce identical output to traditional decoding under greedy sampling.
@@ 7.1 Human Evaluation
where $\mathrm{Win}$ (or $\mathrm{Lose}$) is the number of results generated from the AI search system better (or worse) than the legacy system, and $\mathrm{Tie}$ indicates two results are equally good or bad. In the manual evaluation, we consider the influence of query complexity on the results for the evaluation. Therefore, we sample three categories of simple, moderately complex, and complex queries to comprehensively compare the two systems. Specifically, the test dataset is constructed in advance by sampling real user queries through two distinct selection strategies as follows:
@@ 7.2 Online A/B Test
To evaluate the real-world effectiveness of the AI search system, we deploy it at Baidu Search and conduct an A/B Test comparing it with the legacy search system. When deploying the AI search system at Baidu Search, specific user-side metrics serve as critical indicators of the effectiveness of search systems in enhancing users’ search satisfaction. In the online A/B test, we conduct the experiments with 1.00% real-world web traffic of Baidu search, focusing on metrics directly related to user engagement and reporting the improvements of the proposed AI search system over the online legacy system. As shown in Table 2, compared to the online legacy system, the AI search system achieves the following relative improvements: the change query rate (CQR) decreases by 1.45%; the number of page views (PV) increases by 1.04%; the number of daily active users (DAU) increases by 1.85%; dwell time (Dwell Time) increases by 0.52%. All reported values are statistically significant with p<p<0.05. These results suggest that the proposed AI search system substantially enhances user engagement within the search engine context.